OpenStudy (anonymous):

Find the cube of 20. A. 40 B. 400 C. 800 D. 8000

OpenStudy (anonymous):

If you are talking about volume it would be the side length^3 so 20^3=8000

OpenStudy (whpalmer4):

\[20^3 = 20*20*20\]

OpenStudy (anonymous):

thanks i for got if you had to do that or not thank you both